# Will anyone review this paper? Screening, sorting, and the feedback cycles that imperil peer review

### v2

@@ Abstract
Scholarly publishing relies on peer review to identify the best science. Yet finding willing and qualified reviewers to evaluate manuscripts has become an increasingly challenging task, possibly even threatening the long-term viability of peer review as an institution. What can or should be done to salvage it? Here, we develop mathematical models to reveal the intricate interactions among incentives faced by authors, reviewers, and readers in their endeavors to identify the best science. Two facets are particularly salient. First, peer review partially reveals authors’ private sense of their work’s quality through their decisions of where to send their manuscripts. Second, journals’ reliance on traditionally unpaid and largely unrewarded review labor deprives them of a standard market mechanism—wages—to recruit additional reviewers when review labor is in short supply. We highlight a resulting feedback loop that threatens to overwhelm the peer review system: (1) an increase in submissions overtaxes the pool of suitable peer reviewers; (2) the accuracy of review drops because journals either must either solicit assistance from less qualified reviewers or ask current reviewers to do more; (3) as review accuracy drops, submissions further increase as more authors try their luck at venues that might otherwise be a stretch. We illustrate how this cycle is further propelled by forces including the increasing emphasis on high-impact publications, the proliferation of journals, and competition among these journals for peer reviews. Finally, we suggest interventions that could slow or even reverse this cycle of peer-review meltdown.
@@ I Introduction
When we think about the institution of peer review in science, we often envision reviewers acting akin to shell collectors, sorting through the fragments of shells on a tropical beach in search of whole specimens worth taking home. Indeed, this is part of what peer review does. But papers don’t simply appear on Nature’s editorial desk the way that shells wash up on an expanse of white sand. Authors choose to put them there—and they make those choices in full awareness of the evaluation to follow. Given the stringent review process and the wasted time and effort involved in submitting a paper that is eventually rejected, authors screen their own work, targeting appropriate journals rather than sending everything to the most prestigious outlets. Thus peer review also induces authors to reveal, through their submission decisions, their own private information about the quality of their work [8, 51, 57].
@@
However, this service depends on a massive supply of free labor, namely the unpaid and largely uncredited efforts of peer reviewers [10]. But the pool of peer-review labor has become stretched thin, and the problem seems to be worsening. Editors report increasing difficulty in finding reviewers for the manuscripts that they handle. Bibliometric studies in a range of fields support their assertions: reviewers are more likely to decline review invitations and thus the average number of solicitations per acceptance has increased [[3, 24, 46, 14, 52], though see [58] for an exception]. We appear to be in the midst of a “peer-review meltdown” in which the peer-review system is becoming woefully overtaxed by the volume of manuscript submissions [4, 6, 13, 23, 1].
@@
As peer review teeters, scientists have begun experimenting with new ideas and models to reduce the review load or increase review supply. Examples abound. Platforms such as Publons and Elsevier’s Reviewer Recognition Platform attempt to make reviewing more prestigious by awarding accolades to top reviewers [48, 55]. Brokerage services tried charging authors to secure peer reviews that could be forwarded to prospective publication venues [49, 7, 11]. These foundered, but a new generation of open review platforms such as Review Commons have emerged in their stead [18]. Some computer science conferences such as NeurIPS [44] keep review loads down by disallowing revision and re-review. Elsewhere, some journals offer portable peer review, where initial reviews follow a manuscript if it is resubmitted to other venues. Others have suggested tying the opportunity to submit a paper as an author to one’s contributions as a reviewer [25], and some journals have even experimented with cash payments for reviews [15, 19, 16, 29]. Some critics have even proposed eliminating prepublication peer-review altogether [33]. The breadth of these endeavors testifies to scientists’ eagerness to place scientific publishing on more stable footing.
@@
Yet all these initiatives are hampered by the fact that a rigorous theory of the structure and function of peer review has yet to coalesce [22]. This article aims to begin to fill that gap. While the literature is dotted with mathematical models of peer review, many of these efforts use detailed agent-based models that embrace the richness of the scientific ecosystem (e.g., [50, 5, 36, 38, 20]). In this paper, we take a different approach by developing low-dimensional models that focus on how peer review shifts the burden of identifying the best science among among authors, reviewers, and readers of the scientific literature. (Zhang et al. [56] have recently presented a model of computer-science conferences that also examines how self-screening by authors affects the peer-review burden.) These models reveal a set of hidden pressures on peer review, and bringing them to light helps us to understand the tensions that threaten the entire institution as science changes. In particular, these models highlight a pernicious feedback loop: increasing submissions to top journals exhausts the pool of suitable peer reviewers, resulting in lower quality peer reviews that encourage yet more authors to take a chance on submitting their paper to a prestigious venue (Fig. 1). We also show how other systemic factors—increasing emphasis on high-impact publications, the proliferation of journals, and competition among these journals for unpaid peer-review labor—propel this cycle. Finally, we consider possible solutions that could interrupt the meltdown of peer review, or even reverse it.
@@
Figure 1: The peer-review meltdown cycle. Author screening and journal sorting interact in a feedback loop which inaccurate peer review loosens author screening [2] and looser screening leading to less accurate peer review by depleting the pool of available review labor. Forces that exacerbate this feedback loop include (clockwise from top): Greater rewards to publishing lead more authors to submit their paper to top journals; a proliferation of journals gives authors more opportunities to obtain fresh reviews of already rejected manuscripts; journals’ reliance on a shared pool of review labor compels journals to underuse desk rejection and overexploit the review pool; and noisier review reduces what authors can learn from having their papers rejected, which decreases their selectivity in subsequent submissions.
@@ II.1 Adda-Ottaviani model
Consider a scientific community served by two journals: an elite journal that seeks to publish the top $k\in(0,1)$ manuscripts and a mega-journal that publishes everything else. We consider only two journals for the sake of simplicity, although our model applies in any setting with several vertically differentiated journals. Call $k$ the elite journal’s capacity. We assume that $k$ is determined exogeneously, perhaps by limits imposed by the publisher or by constraints on readers’ attention [26]. Henceforth, we focus on the behavior of the elite journal and place the mega-journal in the background. Thus we refer to the elite journal as simply “the journal”; we say that authors who publish their manuscripts in the elite journal are “published”, and so on.
@@
Suppose that this community contains a unit mass of authors and that each author is endowed with a manuscript with quality $\theta$. For mathematical convenience, assume that $\theta$ has a standard Gaussian distribution across authors, $\theta\sim\mathsf{N}(0,1)$. Authors have their own sense of whether their work is any good. We instantiate this in the model by assuming that each author receives a private signal of their manuscript’s quality. For an author with a manuscript of quality $\theta$, their private signal $X$ is normally distributed with mean $\theta$ and variance $\sigma_{X}^{2}$. Across authors, private signals are marginally distributed as $\mathsf{N}(0,1+\sigma_{X}^{2})$. We refer to the quantile $q=F_{X}(x)$ as the author’s type (here and throughout, $F$ denotes a cumulative distribution function, or cdf), and identify authors with their type, e.g. “author qq”.
@@
Authors can submit their manuscript to the journal, or not. Authors who submit their manuscript pay a nonrecoverable disutility cost $c>0$ [57], which includes the opportunity cost of foregoing immediate publication in the mega-journal. Authors whose manuscripts are published receive kudos, prestige, professional rewards, etc. with value $v>c$. More precisely, $v$ gives the additional reward that the author receives from publishing in the elite journal right away relative to the time-discounted reward of publishing in the mega-journal later. We ignore any other actions that authors may take (e.g., cover letters) that could communicate information about their type to the journal.
@@
Because the journal cannot directly observe manuscript quality $\theta$, it solicits reviews in the usual way. For now, assume that journals send out every manuscript they receive for review; desk rejection will be considered later. Let $Y$ be the review score for a submitted manuscript, and assume that for a manuscript of quality $\theta$, $Y$ is drawn from a Gaussian distribution with mean $\theta$ and variance $\sigma^{2}_{Y}$. Higher review scores $Y$ provide evidence of higher article quality $\theta$, and thus a thresholding rule for $Y$ is rational for a journal that seeks to publish the best manuscripts. Consequently, the journal publishes those papers whose review score exceeds some acceptance threshold $y$.
@@
Two conditions determine the model equilibrium. First, authors submit their paper if and only if their payoff from doing so is positive. We call this the author-rationality condition. [^1] Indifferent authors will comprise a set of measure zero, so their behavior is irrelevant. To be concrete, we assume that indifferent authors submit their manuscript. Second, the journal fills its capacity, a condition we call the capacity-filling condition. The capacity-filling condition can be motivated by assuming that the journal editor prefers to publish as many papers as possible without exceeding the journal’s capacity [41]. (See ref. [56] for a related model in which the elite journal is not capacity-limited but instead seeks to publish all papers with a quality above a particular threshold.)
@@
Figure 2: Graphical illustration of the base model of Adda & Ottaviani [2] and associated welfare measures. A: Equilibrium conditions. The author-rationality condition dictates that the marginal author $\hat{q}$ has acceptance probability $c/v$, and the capacity-filling condition dictates that the area of the gold region—equal to the volume of accepted manuscripts—must equal the journal capacity $k$. Panel based on AO’s Figs. 2 and 3 [2]. For this panel, $k=0.2$, $v/c=5$, and $\sigma_{X}=\sigma_{Y}=1$. B: Welfare measures for the same setting as panel (A). See text for details. C: Change in welfare under greater peer-review noise (in this case, $\sigma_{Y}=2$; all other parameter settings as in panel B). D: Change in welfare as $v/c$ increases (in this case, $v/c=10$; all other parameter settings as in panel B). E: $J=4$ elite journals. F: An infinity of elite microjournals and perfect author knowledge of their manuscript’s quality.
@@ II.2 Welfare
The scientific literature serves (at least) three different constituencies. Authors wish to have their work read. Readers—and by extension, publishers who market journals to those readers—want to read about the best and most interesting science, and do so in a timely fashion. Reviewers contribute the volunteer labor that supports the peer-review process. These constituencies represent roles rather than distinct groups; a given researcher may submit a paper on Monday, read several articles on Tuesday, and write a peer review on Wednesday. But by treating each role as a constituency, we obtain a finer-grained view of how peer review trades off benefits across these activities.
We define payoffs for each group, as follows.
@@ II.2.2 Readers
Readers want to read—and journals want to publish—the best manuscripts. Thus readers’ welfare and the journal’s payoff are the same, and we measure them by the average quality $\theta$ of published manuscripts (which writes as $\mathsf{E}\left[\theta|q\geq\hat{q},Y\geq\hat{y}\right$]) standardized by the average quality of the best $k$ manuscripts ($\mathsf{E}\left[\theta|\theta\geq{F}_{\theta}^{-1}(1-k)\right]$). The readers’ welfare measures how reliably publication indicates scientific quality. This measure of reader welfare is admittedly incomplete, because it ignores the speed with which journals disseminate results; all else equal, readers prefer to learn about new discoveries in a timely fashion. We lack a metric for reader welfare that incorporates timeliness.
@@ II.2.3 Reviewers
While reviewing a manuscript brings a reviewer both benefits and costs, we assume here that the costs to reviewing exceed the benefits. Thus the burden on the review community is proportional to $L$, the review load.
Author and reviewer welfare can be understood graphically (Fig. 2B). The black curve of Fig. 2B gives authors’ benefit (for the same parameter values as Fig. 2A), $v\,\hat{a}(q)$, and the area under this curve, shaded in gray, equals the authors’ total benefit, vkv\,k. The green hatched region has height $c$ and width $L$, and thus its area equals the authors’ total cost, cLc\,L. Author welfare equals the proportion of the gray shaded area that lies outside the hatched green box. The load on reviewers equals the area of the green box divided by $c$.
Before proceeding, we observe that the journal’s publication decisions depend on the combined action of screening and sorting (Fig. 3). The optimal blend of screening and sorting, at least as far as the journal and its readers are concerned, depends authors’ and reviewers’ accuracy in judging manuscripts’ quality. For example, if authors know their manuscript quality’s perfectly ($\sigma_{X}=0$), the journal’s payoff rises as screening becomes more stringent and only the top authors submit their manuscript. Conversely, if reviewers are infallible ($\sigma_{Y}=0$), the journal prefers that all authors submit their manuscript so that the reviewers can find the best ones. The more realistic case in which both authors and reviewers are imperfect judges of manuscript quality intergrades between these two extremes, with the journal’s optimal blend of screening and sorting set to leverage authors’ and reviewers’ knowledge in proportion to the accuracy of each (Fig. A.1).
Figure 3: Author screening and journal sorting combine to make the journal and its readers better off than journal sorting alone. Points show a random sample of 200 manuscripts, located according to the author’s sense of the paper’s strength $X$ and by a reviewer’s score $Y$. Point color indicates the quality of the manuscript $\theta$, with redder (bluer) values corresponding to higher- (lower-) quality manuscripts. The journal publishes $k=20\%$ of the papers, and filled symbols show the actual top 20% of manuscripts. At equilibrium, only authors who think their paper is at least as good as $\hat{x}$ (tan region) submit their paper, and the journal then accepts papers that reviewers rate as at least as good as $\hat{y}$ (intersection of tan and gray regions). The dashed horizontal line shows the acceptance threshold that the journal would use instead if every author submitted their paper. This figure uses $\sigma_{x}=0.5$, $\sigma_{Y}=1$, and $v/c=5$.

### v3

# Screening, sorting, and the feedback cycles that imperil peer review

@@ Abstract
Scholarly journals rely on peer review to identify the science most worthy of publication. Yet finding willing and qualified reviewers to evaluate manuscripts has become an increasingly challenging task, possibly even threatening the long-term viability of peer review as an institution. What can or should be done to salvage it? Here, we develop mathematical models to reveal the intricate interactions among incentives faced by authors, reviewers, and readers in their endeavors to identify the best science. Two facets are particularly salient. First, peer review partially reveals authors’ private sense of their work’s quality through their decisions of where to send their manuscripts. Second, journals’ reliance on traditionally unpaid and largely unrewarded review labor deprives them of a standard market mechanism—wages—to recruit additional reviewers when review labor is in short supply. We highlight a resulting feedback loop that threatens to overwhelm the peer review system: (1) an increase in submissions overtaxes the pool of suitable peer reviewers; (2) the accuracy of review drops because journals either must either solicit assistance from less qualified reviewers or ask current reviewers to do more; (3) as review accuracy drops, submissions further increase as more authors try their luck at venues that might otherwise be a stretch. We illustrate how this cycle is propelled by the increasing emphasis on high-impact publications, the proliferation of journals, and competition among these journals for peer reviews. Finally, we suggest interventions that could slow or even reverse this cycle of peer-review meltdown.
@@ I Introduction
When we think about the institution of peer review in science, we often envision reviewers as acting akin to shell collectors, sorting through the fragments of shells on a tropical beach in search of whole specimens worth taking home. Indeed, this is part of what peer review does. But papers don’t simply appear on Nature’s editorial desk the way that shells wash up on an expanse of white sand. Authors choose to put them there—and they make those choices in anticipation of the evaluation to follow. Given the stringent review process and the wasted time and effort involved in submitting a paper that is eventually rejected, authors screen their own work, targeting appropriate journals rather than sending everything to the most prestigious outlets. Thus peer review also induces authors to reveal, through their submission decisions, their own private information about the quality of their work [1, 2, 3].
@@
However, this service depends on a massive supply of free labor, namely the unpaid and largely uncredited efforts of peer reviewers [4]. But the pool of peer-review labor has become stretched thin, and the problem seems to be worsening. Editors report increasing difficulty in finding reviewers for the manuscripts that they handle. Bibliometric studies in a range of fields support their assertions: reviewers are more likely to decline review invitations and thus the average number of solicitations per acceptance has increased [[5, 6, 7, 8, 9], though see [10] for an exception]. We appear to be in the midst of a “peer-review meltdown” in which the peer-review system is becoming woefully overtaxed by the volume of manuscript submissions [11, 12, 13, 14, 15, 16].
@@
As peer review teeters, scientists have begun experimenting with new ideas to reduce the review load or increase review supply. For example, venues such as Publons and Elsevier’s Reviewer Recognition Platform attempt to make reviewing more prestigious by awarding accolades to top reviewers [17, 18]. Brokerage services tried charging authors to secure peer reviews that could be forwarded to prospective publication outlets [19, 20]. These foundered, but a new generation of journal-independent review initiatives such as Review Commons and Peer Community In have emerged in their stead. Some computer science conferences such as NeurIPS [21] keep review loads down by disallowing revision and re-review. Elsewhere, some journals offer portable peer review, where initial reviews follow a manuscript if it is resubmitted to other venues. Others have suggested tying the opportunity to submit a paper as an author to one’s contributions as a reviewer [22], and some journals have experimented with cash payments for reviews [23, 24, 25, 26]. Yet others have studied whether machine review using AI [27] and large language models [28] can complement peer review. While debate about the propriety of machine review remains unsettled [29, 30], some reviewers are using LLMs for assistance even when journal or conference guidelines forbid doing so [31, 32, 33]. Some critics have even proposed eliminating prepublication peer-review altogether [34]. The breadth of these endeavors testifies to scientists’ eagerness to place scientific publishing on more stable footing.
@@
Yet all these initiatives are hampered by the fact that a rigorous theory of the structure and function of peer review has yet to coalesce [35]. This article aims to begin to fill that gap. While the literature is dotted with mathematical models of peer review, many of these efforts use detailed, agent-based simulation models that embrace the richness of the scientific ecosystem (e.g., [36, 37, 38, 39, 40]). In this paper, we take a different approach by developing low-dimensional models that isolate how peer review shifts the burden of identifying the best science among among authors, reviewers, and readers of the scientific literature. (Zhang et al. [41] have recently presented a model of computer-science conferences that also examines how self-screening by authors affects the peer-review burden.) These models reveal a set of hidden pressures on peer review, and bringing them to light helps us to understand the tensions that threaten the entire institution as science changes. In particular, these models highlight a pernicious feedback loop: increasing submissions to top journals exhausts the pool of suitable peer reviewers, resulting in lower quality peer reviews that encourage yet more authors to take a chance on submitting their paper to a prestigious venue (Fig. 1). We also show how other systemic factors—increasing emphasis on high-impact publications, the proliferation of journals, and competition among these journals for unpaid peer-review labor—propel this cycle. Finally, we consider possible solutions that could interrupt the meltdown of peer review, or even reverse it.
@@
Figure 1: The peer-review meltdown cycle. Author screening and journal sorting interact in a feedback loop which inaccurate sorting loosens author screening [42] and looser screening makes sorting less accurate by depleting the pool of available review labor. Forces that exacerbate this feedback loop include (clockwise from top): Greater rewards to publishing lead more authors to submit their paper to top journals; a proliferation of journals gives authors more opportunities to obtain fresh reviews of already rejected manuscripts; journals’ reliance on a shared pool of review labor compels journals to underuse desk rejection and overexploit the review pool; and noisier review reduces what authors can learn from having their papers rejected.
@@ II.1 Adda-Ottaviani model
Consider a scientific community served by two journals: an elite journal that seeks to publish top manuscripts and a mega-journal that publishes everything else. We consider only two journals for the sake of simplicity, although our model applies in any setting with several vertically differentiated journals. Suppose that the elite journal has the capacity to publish a proportion $k\in(0,1)$ of the manuscripts that the community produces. We assume that the journal’s capacity $k$ is determined exogeneously, perhaps by limits imposed by the publisher or by constraints on readers’ attention [43]. Henceforth, we focus on the behavior of the elite journal and place the mega-journal in the background. Thus we refer to the elite journal as simply “the journal”; we say that authors who publish their manuscripts in the elite journal are “published”, and so on.
@@
Suppose that this community contains a unit mass of authors and that each author is endowed with a manuscript with quality $\theta$. For mathematical convenience, assume that $\theta$ has a standard Gaussian distribution across authors, $\theta\sim\mathsf{N}(0,1)$. Authors have their own sense of whether their work is any good. We instantiate this by assuming that an author with a manuscript of quality $\theta$ obtains a private sense of their manuscript’s quality $X$ that is drawn from a normal distribution with mean $\theta$ and variance $\sigma_{X}^{2}$. Across authors, $X$ is marginally distributed as $\mathsf{N}(0,1+\sigma_{X}^{2})$. We refer to the quantile $q=F_{X}(x)$ as the author’s type (here and throughout, $F$ denotes a cumulative distribution function, or cdf), and identify authors with their type, e.g. “author qq”.
@@
Authors can submit their manuscript to the journal, or not. Authors who submit their manuscript pay a disutility cost $c>0$ [3], which includes the opportunity cost of foregoing immediate publication in the mega-journal. Authors whose manuscripts are published receive kudos, prestige, professional rewards, etc. with value $v>c$. More precisely, $v$ gives the additional reward that the author receives from publishing in the elite journal right away relative to the time-discounted reward of publishing in the mega-journal later. We ignore any other actions that authors may take (e.g., cover letters) that could communicate information about their type to the journal.
@@
Because the journal cannot directly observe manuscript quality $\theta$, it solicits reviews in the usual way. For now, assume that journals send out every manuscript they receive for review; desk rejection will be considered later. Let $Y$ be the review score for a submitted manuscript, and assume that for a manuscript of quality $\theta$, $Y$ is drawn from a Gaussian distribution with mean $\theta$ and variance $\sigma^{2}_{Y}$. Higher review scores $Y$ provide evidence of higher article quality $\theta$, and thus the journal rationally publishes those papers whose review score exceeds some acceptance threshold $y$.
@@
Two conditions determine the model equilibrium. First, authors submit their paper if and only if their payoff from doing so is positive. We call this the author-rationality condition. [^1] Indifferent authors will comprise a set of measure zero, so their behavior is irrelevant. To be concrete, we assume that indifferent authors submit their manuscript. Second, the journal fills its capacity, a condition we call the capacity-filling condition. The capacity-filling condition can be motivated by assuming that the journal editor prefers to publish as many papers as possible without exceeding the journal’s capacity [44]. See ref. [41] for a related model in which the elite journal is not capacity-limited but instead seeks to publish all papers with a quality above a particular threshold.
@@
Figure 2: Graphical illustration of the base model of Adda & Ottaviani [42] and associated welfare measures. A: Equilibrium conditions. The author-rationality condition dictates that the marginal author $\hat{q}$ has acceptance probability $c/v$, and the capacity-filling condition dictates that the area of the gold region—equal to the volume of accepted manuscripts—must equal the journal capacity $k$. Panel based on AO’s Figs. 2 and 3 [42]. For this panel, $k=0.2$, $v/c=5$, and $\sigma_{X}=\sigma_{Y}=1$. B: Welfare measures for the same setting as panel (A). See text for details. C: Change in welfare under greater peer-review noise (in this case, $\sigma_{Y}=2$; all other parameter settings as in panel B). D: Change in welfare as $v/c$ increases (in this case, $v/c=10$; all other parameter settings as in panel B). E: $J=4$ elite journals. Jumps in the cost and benefit curves show the locations of marginal authors $\hat{q}_{i}$, $i=1,\ldots,4$. F: An infinity of elite microjournals and perfect author knowledge of their manuscript’s quality.
@@ II.2 Welfare
The scientific literature serves (at least) three different constituencies. Authors wish to have their work read. Readers—and by extension, publishers who market journals to those readers—want to read about the best and most interesting science, and do so in a timely fashion. Reviewers contribute the volunteer labor that supports the peer-review process. These constituencies represent roles rather than distinct groups; a given researcher may submit a paper on Monday, read several articles on Tuesday, and write a peer review on Wednesday. But by treating each role as a constituency, we obtain a finer-grained view of how peer review trades off benefits across these activities. We define payoffs for each group as follows.
@@ II.2.2 Readers
Readers want to read—and journals want to publish—the best manuscripts. To this extent, readers’ welfare and the journal’s payoff are the same, and we measure them by the average quality $\theta$ of published manuscripts (which can be written as $\mathsf{E}\left[\theta|q\geq\hat{q},Y\geq\hat{y}\right$]) standardized by the average quality of the best $k$ manuscripts ($\mathsf{E}\left[\theta|\theta\geq{F}_{\theta}^{-1}(1-k)\right]$). This measure captures how reliably the journal is able to fill its pages with the best science. As a measure of reader welfare, it is admittedly incomplete because it ignores the speed with which journals disseminate results; all else equal, readers prefer to learn about new discoveries promptly. We lack a metric for reader welfare that incorporates timeliness.
@@ II.2.3 Reviewers
While reviewing a manuscript brings a reviewer both benefits and costs, we assume here that the costs of reviewing exceed the benefits. Thus the burden on the review community scales with $L$, the review load.
Author and reviewer welfare can be understood graphically (Fig. 2B). The black curve of Fig. 2B gives authors’ benefit (for the same parameter values as Fig. 2A), $v\,\hat{a}(q)$, and the area under this curve, shaded in gray, equals the authors’ total benefit, vkv\,k. The green hatched region has height $c$ and width $L$, and thus its area equals the authors’ total cost, cLc\,L. Author welfare equals the proportion of the gray shaded area that lies outside the hatched green box. The load on reviewers equals the area of the green box divided by $c$. Reader and journal welfare depends on the actual article quality $\theta$ and thus is determined by the combined action of screening and sorting (Fig. 3). In the Appendix (Fig. A.1), we show that the optimal strength of screening for the journal and its readers depends on authors’ and reviewers’ accuracy in assessing manuscript quality. These results are intuitive: journals and their readers are better off with more stringent screening and more selective submissions when authors are well-informed about their paper’s quality, and conversely are better off with looser screening and more submissions when referees are better informed.
Figure 3: Author screening and journal sorting combine to determine how capably the journal identifies the most publication-worthy science. Points show a random sample of 200 manuscripts, located according to the author’s sense of the paper’s strength $X$ and by a reviewer’s score $Y$. Point color indicates the quality of the manuscript $\theta$, with redder (bluer) values corresponding to higher- (lower-) quality manuscripts. The journal publishes $k=20\%$ of the papers, and filled symbols show the actual top 20% of manuscripts. At equilibrium, only authors who think their paper is at least as good as $\hat{x}$ (tan region) submit their paper, and the journal then accepts papers that reviewers rate as at least as good as $\hat{y}$ (intersection of tan and gray regions). The dashed horizontal line shows the acceptance threshold that the journal would use instead if every author submitted their paper. This figure uses $\sigma_{x}=0.5$, $\sigma_{Y}=1$, and $v/c=5$.